OpenStudy (magbak):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!! I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!! Can some one explain to me what is sigma notation.

OpenStudy (anonymous):

\[\sum_{}^{} \]

OpenStudy (nory):

Sigma is a sum. Sigma notation asks you to some something. \[\sum_{i=1}^{5} x\] means sum all numbers i from 1 to 5. 1+2+3+4+5 = 15.

OpenStudy (magbak):

The \[\sum_{}^{}\] is sigma notation.

OpenStudy (agent0smith):

It means the sum of, ie the addition of all terms.

OpenStudy (magbak):

Ok and what does the i=1 stand for? Does that refer to imaginary numbers or what does it refer to?

OpenStudy (nory):

It has absolutely nothing to do with imaginary numbers. It just means for i = 1 to i = 5, sum all i's. Like what I said above. (The x is supposed to be an i.)

OpenStudy (magbak):

Ok can you solve one example with showing the step please so I can understand.

OpenStudy (magbak):

Given the arithmetic sequence –6, 1, 8…, evaluate \[\sum_{n=9}^{12}a _{n}\]

OpenStudy (agent0smith):

\[\Large \sum_{k=1}^{4} x^2 \] this means the sum of all x^2's, starting from 1 and ending at 4. It doesn't matter if you use k, or i or whatever letter. \[\Large \sum_{k=1}^{4} x^2= 1^2 + 2^2 +3^2 +4^2 = \]

OpenStudy (magbak):

@Loser66

OpenStudy (agent0smith):

\[\Large \sum_{n=9}^{12}a _{n} = a_9 + a_{10} +a_{11} + a_{12}\] So you need to first find a9, a10, a11, a12. What's the common difference (how much is added from one term to the next) of: –6, 1, 8

OpenStudy (magbak):

You add 7

OpenStudy (magbak):

So what is next.

OpenStudy (agent0smith):

Can you find the general rule? a1 is the first term, d is the difference you just found, n means the n-th term (eg if you want the 9th term, use n = 9). Use this formula to find a9 to a12, then add them up as i showed above. \[\Large a_n = a_1 +(n-1) d\]

OpenStudy (magbak):

What formula.

OpenStudy (magbak):

Oh ok yes I see it now.

OpenStudy (magbak):

But what is a.

OpenStudy (agent0smith):

an means the nth term. a9 means the 9th term, so n=9. Find a9 up to a12.

OpenStudy (magbak):

But how if I do not know what A is.

OpenStudy (agent0smith):

an just means the nth term, whatever term (a number) of the sequence. a is not a variable you need to worry about. You have a1 (the first term) = -6, d=7, so: \[\Large a_n = a_1 +(n-1) d\] becomes \[\Large a_n = -6 +(n-1)*7\] Here is how you find a9 (plug in n = 9: \[\Large a_9 = 6 +(9-1) *7 =???\]

OpenStudy (magbak):

Ok I have another question also if their is a - after the \[\sum -\] does that mean you are subtracting instead of adding.

OpenStudy (magbak):

@agent0smith can you please answer my last question I asked please.

OpenStudy (agent0smith):

Basically, yes, you'd put a negative in front of all the terms (note that a term which is already negative will become a positive).

OpenStudy (magbak):

Ok can we start of baisic like \[\sum_{n=2}^{8}-(-n-5)\]

OpenStudy (agent0smith):

All you have to do is plug in each n, one by one, and add or subtract all the terms: \[\large \sum_{n=2}^{8}-(-n-5) = -(-2-5) -(-3-5) - (-4-5) -...-(-8-5)\] Or to make it easier, you can bring the negative outside the sum: \[\large \sum_{n=2}^{8}-(-n-5) = - \large \sum_{n=2}^{8}(-n-5) \] or just distribute the negative into the brackets \[\large \sum_{n=2}^{8} (n+5) = (2+5) + (3+5) + ... + (8+5)\]

OpenStudy (magbak):

Wait but how did n become 3.

OpenStudy (agent0smith):

By plugging in n=2, n=3, n=4... and so on.

OpenStudy (agent0smith):

That sum means start from n=2, and stop at n=8. All you have to do is plug in each n.

OpenStudy (magbak):

Oh so you just skip 1.

OpenStudy (magbak):

So the answer is 70

OpenStudy (agent0smith):

You start at whatever the bottom number is, in this case the n=2.

OpenStudy (magbak):

Yes I mean after you add everthing up and the final answer is 70. can you just check my answer please.

OpenStudy (magbak):

Yes I am correct.

OpenStudy (magbak):

Thank you so much.

OpenStudy (agent0smith):

You're welcome.

OpenStudy (magbak):

@agent0smith

OpenStudy (magbak):

@agent0smith

OpenStudy (agent0smith):

yes...?

OpenStudy (magbak):

Oh I was trying to find the answer to the first question that I found hard.

OpenStudy (magbak):

But I got it thank you.